A die is thrown once what probability that it is a six?
It is 1/6. (It is just the same if the favorable outcome is 1,2,3,4 or 5, provided that it is a fair dice.)
A friend used this phrase. I think she meant Boolean Algebra.
What is the probability of choosing a red ten?
Since there are 2 red tens in a deck of 52 cards, the probability of choosing a red ten is 2 out of 52 = 1 out of 26 = 0,0384615385...
What phenotype has the greatest frequency in a trait that follows a normal distribution?
a phenotype with the greatest frequency in a trait is enviromental conditions. (G00D !-U<K)
A die is rolled 10 times what the probbility all are 6?
The probability of rolling a 6 is 1/6. The probability of rolling 10 times a 6 is (1/6)10 or 1.654X10-8.
3/7*3/7= 9/49
Should studens be allowed to ride bike in school?
No because it is dangrous and you might get heart you might get in trouble
What are the odds of rolling a 1 in dice?
The probability of rolling a 1 on a die is 1/6 if you roll it once.
68% of the scores are within 1 standard deviation of the mean -80, 120
95% of the scores are within 2 standard deviations of the mean -60, 140
99.7% of the scores are within 3 standard deviations of the mean -40, 180
The probability of ac and bc is 1/5.
How many ways can you arrange 5 people for a photo?
I think this is a homework question so I'm giving only the method.
The answer is 5-factorial, expressed as 5!,
where the factorial of a number N is the product of:
N X (N-1) X (N-2) X .... X 1
E.g. 3! = 3 X 2 X 1 = 6.
What is the probability that in four children two will be feeble-minded girls?
It is not possible to answer the question for a number of reasons.
The gender of children are not independent of one another. They depends on the parents.
Second, the feeble-mindedness of the child or otherwise is greatly influenced by their nurture: the family and school play an important role in that context. This will affect feeble-minded boys equally.
Exactly what it says.
The precise location of an electron cannot be ascertained. The best that can be done is to calculate probabilities that it is found within various locations.
How many ways can Zoe choose 4 figurines from a collection of 7?
(7 x 6 x 5 x 4) = 840 different ways.
But each (4 x 3 x 2) = 24 ways result in the same setof 4 .
So there are only 840/24 = 35 different sets of 4 figurines.
How many times do the dice roll on odd numbers?
As many times as you like. If there are not enough times, keep on rolling them, get your family and friends to join you.
What does probability look like?
Probability is an abstract concept and so does not have any particular appearance.
What is the probability of pulling a green or yellow marble from the bag?
It depends on whether their is those colors of marbles in the bag, and how many of each, and if so you have to put it into ratios or percents and the one with the larger number is the one with the greater chance... sincerely,
Grade 8 Student....in Canada
There are 26 red cards and 26 black cards.
3 red cards can be chosen in 26C3 ways
2 black cards can be chosen in 26C2 ways
The required answer is 26C3 X 26C2 ways.
Answer: 1067742
S Suneja
The data from a normal distribution are symmetric about its mean, not about zero. There is, therefore nothing strange about all the values being negative.
The probability is 21/36 = 7/12
What is the probability of a couple having a girl?
It depends on the couples' genes. Also, at present the probability of a girls is approx 0.48